sin / cos =?

sin / cos =?

Grade:12

14 Answers

venky
8 years ago
tan
KARTHIK
8 years ago
TAN
SHANMUKESHWAR
8 years ago
tan
chandu
8 years ago
tan
chandu
8 years ago
tan
chandu
8 years ago
tan
chandu
8 years ago
tan
gopathoti bujji
8 years ago
tan
Purna Pavan Chandra
8 years ago
TAN
SAPTAGIRI
8 years ago
tan
Mondo Camry
8 years ago
we cant say anything bcoz you have not mentioned any angle
VENKATA SAI
8 years ago
tan
Rahul Jiji George
8 years ago
sin theta  /cos theta = P/H / A/H = P/A =TAN theta  this is for any acute angle i.e lies in the firstquadrant
Forum Team
8 years ago
Dear Students,
 
We noticed that some students are copying and pasting same answer that has posted by previous users. Also, they are creating multiple account to post question and then answering those questions & approving as well. Please note that all these activities come under spamming and no gift voucher will be issued to them.
 
Hope that you will maintain the decorum of this forum and help the students genuinely
 
Forum Team